### v1

# RiemannFormer: A Framework for Attention in Curved Spaces

Abstract
This research endeavors to offer insights into unlocking the further potential of transformer-based architectures. One of the primary motivations is to offer a geometric interpretation for the attention mechanism in transformers. In our framework, the attention mainly involves metric tensors, tangent spaces, inner product, and how they relate to each other. These quantities and structures at discrete positions are intricately interconnected via the parallel transport of tangent vectors. To make the learning process more efficient, we reduce the number of parameters through ingenious predefined configurations. Moreover, we introduce an explicit mechanism to highlight a neighborhood by attenuating the remote values, given that transformers inherently neglect local inductive bias. Experimental results demonstrate that our modules deliver significant performance improvements relative to the baseline. More evaluation experiments on visual and large language models will be launched successively.
Keywords Transformer  $\cdot$ Self-Attention  $\cdot$ Curved Space  $\cdot$ Tangent Mapping  $\cdot$ Parallel Transport
1 Introduction
The transformer architecture, introduced in the seminal paper by Vaswani et al. [1], has revolutionized the field of natural language processing (NLP) and extended its impact to computer vision, audio processing, and other domains. The iterative application of transformer blocks allows tokens to progressively capture higher-level semantic concepts, making the architecture highly effective for tasks requiring contextual understanding.
The transformer block operates on a sequence of tokens, each represented as a vector of fixed dimensionality $d_{\text{model}}$. For a sequence of $L$ tokens, the input to the block is a matrix $X\in\mathbb{R}^{L\times d_{\text{model}}}$, where each row corresponds to a token’s embedding. The transformer block processes this matrix through its core component, Multi-Head Self-Attention (MHSA). The MHSA mechanism is the cornerstone of the transformer, enabling it to aggregate information globally across all tokens in the input sequence. Unlike convolutional neural networks (CNNs), which rely on local receptive fields, or recurrent neural networks (RNNs), which process sequences sequentially, self-attention allows each token to attend to all other tokens simultaneously, capturing long-range dependencies efficiently. Specifically, the attention mechanism computes a score for each pair of tokens by taking the dot product of their query and key vectors, scaled to prevent large values that could destabilize training: $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}$. The dot product $\mathbf{QK}^{\top}\in\mathbb{R}^{L\times L}$ measures the similarity between the query vector of one token and the key vectors of all tokens, producing an attention score matrix. The softmax function normalizes the scores across the key dimension, converting them into weights that sum to $1$ for each query. This ensures that the attention mechanism focuses on the most relevant tokens. The resulting weights are used to compute a weighted sum of the value vectors, producing a new representation for each token that incorporates information from all other tokens. The output of the attention mechanism is a matrix, where each row is a context-aware representation of the corresponding input token. Moreover, the attention mechanism is permutation-equivariant, meaning the output does not depend on the order of the input tokens unless positional information is explicitly added.
Since the self-attention mechanism is permutation-equivariant, it does not inherently capture the order of tokens. To address this, transformers incorporate positional encodings or positional embeddings, which are added to the input token embeddings before processing. Below, we provide a brief overview of positional encodings in transformer. Positional encodings are indispensable in transformer models, embedding sequential order into permutation-equivariant self-attention mechanisms to enable token position modeling in tasks like natural language processing and vision. The seminal sinusoidal positional encodings, introduced by Vaswani et al. [1], remain prevalent due to their deterministic nature and ability to generalize across arbitrary sequence lengths. In contrast, learned positional encodings, as used in BERT [2], optimize trainable embeddings for task-specific patterns but falter in extrapolating to longer sequences. Relative positional encodings, adopted in Transformer-XL [3], T5 [4], and TUPE [5], model token distance relationships, enhancing generalization for long-context tasks. Rotary Positional Embeddings (RoPE) efficiently encode relative positions by applying rotation matrices to query and key vectors [6], with extensions like xPos [7] incorporating dynamic scaling for stable long-sequence training. Recent innovations include ALiBi [8], which simplifies computation by introducing attention biases based on token distances. Prior methods like absolute and relative positional encoding are static post-training. Zheng et al. propose Data-Adaptive Positional Encoding (DAPE), which adjusts dynamically based on input context and learned priors [9]. In addition, Golovneva et al. introduced Contextual Position Encoding (CoPE) [10], a novel position encoding method that conditions positions on context by selectively incrementing positions based on model-determined tokens. Although methods like ALiBi and adaptive PEs reduce complexity, they may introduce instability in extremely long sequences, and their generalization across diverse domains remains underexplored.
Since its adaptation to visual tasks with the Vision Transformer (ViT) [11], the transformer has demonstrated exceptional performance across a wide range of applications, from high-level tasks like classification [12] and detection [13] to low-level tasks such as super-resolution and generation. ViTs process images by dividing them into fixed-size patches, which are then treated as tokens similar to words in a sentence. These patches are embedded and fed into a transformer encoder, which leverages self-attention mechanisms to capture global relationships within the image. This approach allows ViTs to learn complex patterns and dependencies across the entire image, leading to impressive performance on various benchmarks. Positional encoding is a crucial component in the application of transformers to visual tasks, enabling the model to retain spatial information about the input image. Unlike natural language processing tasks where the sequence of words carries inherent positional information, images are divided into patches that lack such sequential context. Positional encodings allows the model to account for the order and arrangement of the patches, which is essential for understanding the spatial relationships and structure within the image. The original ViT model uses fixed positional encodings, similar to those used in the original Transformer model for NLP tasks. In addition to learnable and relative positional encodings, alternative schemes have emerged. RoPE excels in language models, particularly in enhancing transformer length extrapolation, yet its potential in computer vision remains largely unexplored. Heo et al. analyzes RoPE’s application to ViT, demonstrating significant improvements in visual tasks like classification, detection and segmentation, with minimal computational overhead [14]. Chu et al. introduce conditional positional encoding for ViT, dynamically generated based on local input token context [15]. These approaches potentially provide a more flexible and scalable approach to positional encoding in ViTs.
Despite its strengths, positional encodings require further interpretation and improvement. The mechanistic role of positional information remains unclear, offering a promising direction for future research. To better elucidate the attention mechanism, we endeavor to investigate the deeper implications underlying positional encoding in this paper. In addition, transformers often benefit from being trained on massive datasets. Larger datasets allow transformers to learn more nuanced patterns and relationships in the data, leading to better accuracy and generalization. However, transformers often exhibit suboptimal performance when applied to small datasets. Through experimental analysis, we have found that the introduction of an explicit locality focusing mechanism significantly enhances the performance of transformer in visual tasks.
2 Methodology
Consider a sequence comprising $L$ input tokens, with their embeddings represented as $\{\boldsymbol{x}_{i}\}_{i=1}^{L}$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{D}$ denotes the $D$-dimensional embedding vector for the $i$-th token, devoid of positional information. In this paper, we adopt an alternative perspective: we posit the existence of a curved space $\mathcal{M}$ wherein the encoding of a token sequence corresponds to identifying a sequence of points $\{\mathbf{p}_{i}\}$, or a grid of points $\{\mathbf{p}_{ij}\}$ for image data.
The geometric structure of a curved space implicitly supports positional information for sequences or grids of points. However, in such a non-flat space, vectors in tangent spaces at distinct points cannot be directly operated upon. What strategies can be employed to effectively tackle this challenge? Specifically, the vectors $\mathbf{q}_{i}$ and $\mathbf{k}_{i}$ reside in the tangent spaces $T_{\mathbf{p}_{i}}\mathcal{M}$ at position $\mathbf{p}_{i}$. For instance, $\mathbf{q}_{m}$ and $\mathbf{k}_{m}$ both belong to the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$, whereas $\mathbf{q}_{m}$ and $\mathbf{k}_{n}$ are situated in the tangent spaces at points $\mathbf{p}_{m}$ and $\mathbf{p}_{n}$, respectively. The inner product in the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$ is defined as $\mathbf{x}_{1}^{\top}\mathbf{M}_{m}\mathbf{x}_{2}$, where $\mathbf{M}_{m}$ is a Riemannian metric specifying the inner product structure at $\mathbf{p}_{m}$. Since $\mathbf{k}_{n}$ does not belong to the tangent space at $\mathbf{p}_{m}$, it is not appropriate to directly perform inner products such as $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathbf{k}_{n}$. To address this, a transformation $\mathtt{f}_{n,m}(\mathbf{k}_{n})$ is required to appropriately transport $\mathbf{q}_{n}$ from the tangent space at $\mathbf{p}_{n}$ to that at $\mathbf{p}_{m}$. Only after this transformation can a valid inner product be executed, given by $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathtt{f}_{n,m}(\mathbf{k}_{n})$.
In Euclidean space, tangent spaces at various points are isomorphic via translation, enabling straightforward vector transfer. This identification stems from the flat, homogeneous nature of Euclidean geometry, allowing parallel vector movement without additional geometric considerations. On a Riemannian manifold, parallel transport moves vectors along curves, maintaining consistency with intrinsic geometry structure of the manifold during transfer between tangent spaces. To be more specific, the vectors $\mathbf{q}_{m}$ and $\mathbf{k}_{n}$ ($m\neq n$) are in tangent spaces at different positions $\mathbf{p}_{i}$ respectively. Prior to performing the inner product, it is necessary to adapt the vectors to the same space. This can be accomplished, for instance, by applying a parallel transport, denoted as $\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})$, that transports $\mathbf{k}_{n}$ into the tangent space at $\mathbf{q}_{m}$.
Preliminaries: The parallel transport of tangent vectors $\mathbf{v}\in T_{c(\alpha)}\mathcal{M}$ to the tangent space $T_{c(\beta)}\mathcal{M}$ along a curve $c(t):[0,1]\to\mathcal{M}$ is denoted by $\Gamma_{c[\alpha,\beta]}(\mathbf{v})$. Two fundamental properties can be leveraged:
1) $\Gamma_{c[\alpha,\beta]}:\;T_{c(\alpha)}\mathcal{M}\to T_{c(\beta)}\mathcal{M}$ is a linear map.
2) Consider a Riemannian manifold $\mathcal{M}$ equipped with its associated Riemannian connection $\nabla$. Let $\mathbf{V}$ be a parallel vector field along a parameterized curve $c(t):[0,1]\to\mathcal{M}$. The vector field $\mathbf{V}$ maintains a constant magnitude along the curve $c$.
To summarize, parallel transport provides a way of moving tangent vectors along a curve using the affine connection, and this provides a linear isomorphism between the tangent spaces at the two ends of the curve. However, the isomorphism obtained in this way will in general depend on the choice of the curve. Furthermore, it must be noted that we lack knowledge of the positions of points, curves, and additional structures (such as explicitly defined connection). Consequently, how can we achieve the parallel transport of tangent vectors? In fact, our methodology does not involve the explicit positional embeddings. Instead, we appoint the points and the curves connecting them in the order of tokens and enable the model to learn the linear mappings between tangent spaces at different points. Through these linear mappings, we can perform valid inner product without involving specific positional information.
Leveraging the second property mentioned above, we derive the following equation:
$\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}=\mathcal{P}_{n\rightarrow m}({\bf k}_{n})^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}({\bf k}_{n})$ (1)
First, we introduce a common reference space $\mathbb{R}^{D}$, upon which the linear transformations between tangent spaces at various points are subsequently learned. The linear transformation to be learned at point $\mathbf{p}_{i}$ is denoted by $\mathcal{P}_{i}:\mathbb{R}^{D}\to T_{\mathbf{p}_{i}}\mathcal{M}$ and can be represented by a $D\times D$ matrix $\mathbf{T}_{i}$. In this study, we term the transformation $\mathcal{P}_{i}^{-1}$ as tangent mapping or tangent aligning.
Building upon the foregoing analysis, we have
$\mathcal{P}_{n\rightarrow m}({\bf k}_{n})=\mathcal{P}_{m}\circ\mathcal{P}^{-1}_{n}({\bf k}_{n})=\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}.$ (2)
By substituting this result into Eq. (1), the following equation can be derived:
$\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}=\mathbf{k}_{n}^{\top}\left(\mathbf{T}^{-1}_{n}\right)^{\top}\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}$ (3)
Given the arbitrariness of $\mathbf{k}_{n}$, it consequently follows that:
$\mathbf{M}_{n}=\left(\mathbf{T}^{-1}_{n}\right)^{\top}\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}.$ (4)
Then, simultaneously left-multiplying by matrix $\mathbf{T}_{n}^{\top}$ and right-multiplying by matrix $\mathbf{T}_{n}$, we obtain:
$\mathbf{T}_{n}^{\top}\mathbf{M}_{n}\mathbf{T}_{n}=\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}.$ (5)
The above relationship of these matrices is what we must consider when predefining them.
To clarify again, the aim of this work is to enable the model to learn the geometric structure of the curved space where the sequence data reside. The primary determinants are the metric tensors and the transformations between the tangent spaces at the positions involved. Given the intricate nature and the extensive array of parameters involved, it is essential to establish predefined configurations for these matrices.
A Riemannian metric $\mathbf{M}_{i}$ can be represented by a symmetric positive definite matrix, but for the sake of simplifying the problem, we preset it as a diagonal matrix:
$\mathbf{M}_{m}=\begin{bmatrix}s_{1}^{m}&&&\\ &s_{2}^{m}&&\\ &&\ddots&\\ &&&s_{D}^{m}\end{bmatrix},$ (6)
where $s_{i}>0$ are learnable parameters. To further reduce the number of parameters, a simpler presetting, a scalar matrix $\mathbf{M}_{m}=s^{m}\mathbf{I}$ can be employed, where $\mathbf{I}$ is the identity matrix. This is also known as a conformal deformation of $\mathbb{R}^{D}$. To impose the constraint $s<1$, we may proceed by defining it as $s=\frac{\exp(w)}{\exp(w)+\alpha}$, where $w$ is a learnable parameter and $\alpha$ denotes a hyperparameter. Alternatively, one might opt for $s=\exp(w)$, which permits $s>1$. For long sequences, we can decelerate the rate of exponential growth, by setting the parameter $s^{\log_{\beta}(m)}$. In addition, the learnable metric matrix of general form is given in the Appendix section.
To facilitate the model to effectively learn these linear transformations, we first decompose them into two components, including scaling and length-preserving in $\mathbb{R}^{D}$. The introduction of scaling part is motivated by the characteristics of the curved space and the existence of Riemannian metric. The scaling part is represented by a diagonal matrix or a scalar matrix $s\mathbf{I}$. An orthogonal matrix is used for length-preserving, and a rotation is naturally a candidate. A rotation in $n$-dimensional Euclidean space is a linear transformation that preserves lengths of vectors, and it can be represented by an $D\times D$ orthogonal matrix in the special orthogonal group $SO(D)$. It is well known that an element of $SO(D)$ can be generated from the exponential of a skew-symmetric matrix $\mathbf{X}$, denoted as $\exp(\mathbf{X})$.
For position $i$, we predefine the linear transformation as,
$\mathbf{T}_{i}=s^{-\frac{i}{2}}\exp(i\mathbf{X}),$ (7)
Subsequently, we obtain,
$\mathbf{T}_{m}\mathbf{T}_{n}^{-1}$ $=s^{-\frac{m}{2}}\exp(m\mathbf{X})\cdot s^{\frac{n}{2}}\exp(-n\mathbf{X})$ (8)
$=s^{\frac{n-m}{2}}\exp((m-n)\mathbf{X})$
and
$\langle\mathbf{q}_{m},\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})\rangle_{\mathbf{M}_{m}}$ $=\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$ (9)
$=\mathbf{q}_{m}^{\top}\left[s^{m}s^{\frac{n-m}{2}}\exp((m-n)\mathbf{X})\right]\mathbf{k}_{n}$
$=\mathbf{q}_{m}^{\top}\left[s^{\frac{m+n}{2}}\exp((m-n)\mathbf{X})\right]\mathbf{k}_{n}$
$=\langle s^{\frac{m}{2}}\exp(-m\mathbf{X})\mathbf{q}_{m},\;s^{\frac{n}{2}}\exp(-n\mathbf{X})\mathbf{k}_{n}\rangle_{\mathbf{I}}$
$=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$
It should be noted that a skew-symmetric matrix has ${D(D-1)}/{2}$ independent entries and thereby quite a few learnable parameters will be introduced. Similarly, in order to reduce the number of learnable parameters, We can also formally predefine them as block diagonal matrices.
In this work, the notation $\breve{\mathbf{R}}(\theta)$ is used to represent a counterclockwise rotation through an angle $\theta$ in a 2D subspace. We predefine the transformation matrices as follows,
$\mathbf{T}_{m}=s^{-\frac{m}{2}}\begin{bmatrix}\breve{\bf R}(m\theta_{1})&&&\\ &\breve{\bf R}(m\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}(m\theta_{D/2})\end{bmatrix},$ (10)
and the inverse matrices,
$\mathbf{T}_{m}^{-1}=s^{\frac{m}{2}}\begin{bmatrix}\breve{\bf R}(-m\theta_{1})&&&\\ &\breve{\bf R}(-m\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}(-m\theta_{D/2})\end{bmatrix}.$ (11)
Then, the product is
$\mathbf{T}_{m}\mathbf{T}_{n}^{-1}=s^{\frac{n-m}{2}}\begin{bmatrix}\breve{\bf R}((m-n)\theta_{1})&&&\\ &\breve{\bf R}((m-n)\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}((m-n)\theta_{D/2})\end{bmatrix}.$ (12)
By substituting it into Eq. (1), we have
$\langle\mathbf{q}_{m},\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})\rangle_{\mathbf{M}_{m}}=\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})=\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\mathbf{k}_{n}$ (13)
$={\bf q}_{m}^{\top}s^{\frac{n+m}{2}}\begin{bmatrix}\breve{\bf R}((m-n)\theta_{1})&&&\\ &\breve{\bf R}((m-n)\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}((m-n)\theta_{D/2})\end{bmatrix}\mathbf{k}_{n}.$
It is important to note that the factor $s^{\frac{n+m}{2}}$ disrupts the relativity concerning the positions if $s\neq 1$. Such a mechanism also endows our framework with the capability to balance the capture of both absolute and relative position information. The model autonomously emphasizes the relative information during the learning process by adjusting the parameter $s$. Consequently, this relationship underscores the versatility and broader applicability of our framework. It is also worth noting that, under the aforementioned configuration, the proposed approach reduces to a conventional RoPE when setting $s=1$, thereby indicating that RoPE can be viewed as a special case of our framework.
Furthermore, Eq. (13) can be transformed into the form of a standard inner product,
$\langle s^{\frac{m}{2}}\breve{\bf R}(-m\theta)\mathbf{q}_{m},\;s^{\frac{n}{2}}\breve{\bf R}(-n\theta)\mathbf{k}_{n}\rangle_{\mathbf{I}}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$ (14)
Note that the above formulation represents a simplified version involving only a 2D subspace.
In addition, when considering the transformation matrix $\mathbf{T}_{i}$, we can also use $2\times 2$ reflection matrices as blocks. The notation $\hat{\mathbf{R}}(\theta)$ is used to represent a reflection matrix
$\begin{bmatrix}\cos(2\theta)&\;\;\;\;\!\sin(2\theta)\\ \sin(2\theta)&-\cos(2\theta)\end{bmatrix}.$ (15)
This matrix represents the reflection of an arbitrary two-dimensional vector in a 2D subspace about the axis defined by the vector $\mathbf{d}=(\cos(\theta),\sin(\theta))$. The determinant of a reflection matrix $\hat{\mathbf{R}}(\theta)$ is $-1$, indicating that it reverses orientation. The product of two reflections returned to the original orientation again. Furthermore, this reflection matrix satisfies the property $\hat{\mathbf{R}}(\theta)^{-1}=\hat{\mathbf{R}}(\theta)$, indicating that the matrix is its own inverse. Moreover, the product of rotation matrices and reflection matrices $\widetilde{\bf R}(\theta_{1},\theta_{2})=\hat{\bf R}(\theta_{2})\breve{\bf R}(\theta_{1})$ are still reflection matrices, and the product of two reflection matrices yields a rotation matrix: $\breve{\bf R}(\theta_{1}-\theta_{2})=\hat{\bf R}(\frac{\theta_{1}}{2})\hat{\bf R}(\frac{\theta_{2}}{2})$.
Consequently, beyond rotation matrices, reflection matrices can also be leveraged for the predefinition of the transformation matrix. For example, the rotation matrices in Eq. (10) may be directly substituted as reflection matrices as follows,
$\hat{\mathbf{T}}_{i}=s^{-\frac{i}{2}}\begin{bmatrix}\hat{\mathbf{R}}(i\theta_{1})&&&\\ &\hat{\mathbf{R}}(i\theta_{2})&&\\ &&\ddots&\\ &&&\hat{\mathbf{R}}(i\theta_{D/2})\end{bmatrix}.$ (16)
The subsequent operations follow a similar procedure as previously described and, for the sake of brevity, are omitted here.
We can also predefine the transformations in higher-dimensional subspaces by combining reflections and rotations. Such block matrices as below,
$\hat{\mathbf{R}}(\theta_{1},\theta_{2})=\begin{bmatrix}\breve{\bf R}(\theta_{1})&\mathbf{0}\\[6.79993pt] \mathbf{0}&\hat{\bf R}(\theta_{2})\end{bmatrix}_{4\times 4},$ (17)
and their inverse matrices are
$\hat{\mathbf{R}}(\theta_{1},\theta_{2})^{-1}=\begin{bmatrix}\breve{\bf R}(-\theta_{1})&\mathbf{0}\\[6.79993pt] \mathbf{0}&\hat{\bf R}(\theta_{2})\end{bmatrix}_{4\times 4}.$ (18)
Such combined block matrices may further broaden the space of predefined transformation matrices and allow the model to explore more complex geometric manipulations and representations, thereby gaining deeper insights into the underlying structures and relationships within the data.
2.1 Locality Focusing
Unlike CNNs with locality and translational equivariance, vanilla transformers are data-hungry due to their minimal inductive biases. This flexibility makes them powerful on large, diverse datasets but fragile on limited data, where domain-specific priors are crucial for generalization. To enhance the local inductive bias of transformers, we introduce an explicit mechanism to highlight a neighborhood by attenuating the remote values.
Assume that the tokens are placed in a flat space and the explicit positional embedding are also absent, is it possible to introduce an additional mechanism to address the lack of awareness of sequence order in self-attention module?
To clarify, rather than considering the values $\mathbf{V}$ in a curved space, we aggregate them in the parameter space $\mathbb{R}^{D}$. Given a position $m$, we introduce a decay factor for the values $\mathbf{V}$ to attenuate the influence of remote positions. Specifically, we introduce factors $\lambda_{i}$ to scale the values $\mathbf{v}_{i}$ and update the attention function $\sum_{i}s_{i}(\lambda_{i}\mathbf{v}_{i})$. We compute the output vector as:
$\mathbf{v}_{m}^{\prime}(\mathbf{q}_{m})=\text{Attention}(\mathbf{q}_{m},\mathbf{K},\mathbf{\Lambda},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{q}_{m}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)(\mathbf{\Lambda}\mathbf{V})=\mathbf{s}(\mathbf{\Lambda}\mathbf{V}),$ (19)
where $\mathbf{q}_{m},\;\mathbf{s}\in\mathbb{R}^{1\times D}$, $\mathbf{K},\;\mathbf{V}\in\mathbb{R}^{L\times D}$, and a diagonal matrix $\mathbf{\Lambda}\in\mathbb{R}^{L\times L}$. The formula can be expanded as follows,
$\mathbf{v}_{m}^{\prime}(\mathbf{q}_{m})$ $=\begin{bmatrix}s_{m1}&s_{m2}&\cdots&s_{mL}\end{bmatrix}\begin{bmatrix}\lambda_{m1}&&&\\ &\lambda_{m2}&&\\ &&\ddots&\\ &&&\lambda_{mL}\end{bmatrix}\begin{bmatrix}\mathbf{v}_{1}\\[3.6pt] \mathbf{v}_{2}\\ \vdots\\ \mathbf{v}_{L}\end{bmatrix}$ (20)
$=\begin{bmatrix}s_{m1}&s_{m2}&\cdots&s_{mL}\end{bmatrix}\begin{bmatrix}\lambda_{m1}\mathbf{v}_{1}\\[3.6pt] \lambda_{m2}\mathbf{v}_{2}\\ \vdots\\ \lambda_{mL}\mathbf{v}_{L}\end{bmatrix}.$
Here, we discuss $2D$ image data, although this mechanism is equally applicable to $nD$ sequence data. The output value $\mathbf{v}^{\prime}_{m}$ can also be interpreted from the perspective of a bilateral filter [16] in the context of image processing,
$\mathbf{v}^{\prime}_{m}=\sum_{n=1}^{L}{G}_{s}(\mathbf{q}_{m},\mathbf{k}_{n})G_{\sigma}(\mathbf{p}_{m},\mathbf{p}_{n})\mathbf{v}_{n},$ (21)
where $G_{s}(\cdot,\cdot)$ indicates the scaled dot-product of the query and key vectors followed a softmax function, $\mathbf{p}_{m}$ and $\mathbf{p}_{n}$ are the $2D$ positions of image patches on the image grid, $G_{\sigma}(\cdot,\cdot)$ is an attenuation factor calculated based on the norm of the difference between $\mathbf{x}$ and $\mathbf{y}$, such as $G_{\sigma}(\mathbf{x},\mathbf{y})=e^{-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}_{A}}{2\sigma^{2}}}$. $\mathbf{A}$ is a symmetric positive-definite matrix such that $\|\mathbf{v}\|_{\bf A}^{2}=\mathbf{v}^{\top}\mathbf{A}\mathbf{v}$, it is learnable and is set to the identity matrix by default. $\sigma$ is a learnable parameter that can be assigned a unique value for each position or shared as one parameter for all positions in an attention head. From the perspective of the bilateral filter, $G_{s}(\cdot,\cdot)$ controls intensity similarity and $G_{\sigma}(\cdot,\cdot)$ controls spatial attenuating.
In the implementation, the attenuation factors corresponding to one position are reshaped into a row, thus the ones of all positions are packed together into a $L\times L$ attenuation matrix. For all tokens, we can first perform element-level multiplication on the score matrix and the attenuation matrix $\mathbf{\Omega}$, and then perform matrix multiplication on the result matrix and $\mathbf{V}$. We have
$(\mathbf{S}\odot\mathbf{\Omega})\cdot\mathbf{V}=\left(\begin{bmatrix}s_{11}&s_{12}&\cdots&s_{1L}\\ s_{21}&s_{22}&\cdots&s_{2L}\\ \vdots&\vdots&\ddots&\vdots\\ s_{L1}&s_{L2}&\cdots&s_{LL}\end{bmatrix}\odot\begin{bmatrix}\lambda_{11}&\lambda_{12}&\cdots&\lambda_{1L}\\ \lambda_{21}&\lambda_{22}&\cdots&\lambda_{2L}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{L1}&\lambda_{L2}&\cdots&\lambda_{LL}\end{bmatrix}\right)\cdot\mathbf{V}.$ (22)
Such a simple mechanism is introduced to improve the capacity of transformers to capture and learn the inductive bias of locality especially on small datasets. Our experimental results reveal that incorporating such an explicit local focusing mechanism substantially improves the performance of transformers in visual tasks.
3 Experiments
This study primarily focuses on the visual domain and, consequently, evaluates the performance of RiemnnFormer on image classification tasks for small datasets.
In the context of image data, there are two axes to consider. Although these axes can be addressed jointly [14] or the image patches are directly flatten into a sequence, we are interested in exploring the effects of processing them separately along the $x$ and $y$ axes. This strategy also facilitates a comparative studies with previous work.
The proposed modules undergo training and evaluation on the CIFAR-10 and CIFAR-100 datasets, employing a lightweight Vision Transformer (ViT) architecture. Specifically, these modules leverage the compact yet efficient structure of the ViT model to process the image classification tasks inherent to these benchmark datasets.
All experiments were implemented on an NVIDIA RTX 4090 GPU with PyTorch 2.7. The training procedure is conducted over 200 epochs with a batch size of 128, utilizing the AdamW optimizer configured with its default hyperparameters.
In predefined transformations, certain operations (such as rotation $\breve{\bf R}(\theta)$ and reflection $\hat{\bf R}(\theta)$) involve the configuration of angles. An angle schedule is a function that sends an integer to an angle involved in the linear transformation. The choice of the angle schedule can also impact the performance of the tangent mapping. In our experiments, we observed that the following straightforward schedule yields favorable results: $f(t):t\mapsto t\theta,\;t\in\{1,2,\dots\}$, where $\theta$ is a learnable parameter or a predefined angle. Exploring more effective angle schedules is also a direction worth investigating. In the current experiment of this work, we use the predefined configuration in Eq. (10). We set the hyperparameter $\alpha=0.1$ in $s=\frac{\exp(w)}{\exp(w)+\alpha}$ in these experiments.
The current work is primarily dedicated to further exploring the potential of transformer-based models in visual tasks, and thus, it mainly compares with ViT architectures. The architecture of our ultra-lightweight vision transformer comprises only $4$ attention heads, $6$ transformer layers, and decomposes images into $64$ patches, each of $4\times 4$ pixels. The results from the experiments are compiled in Table 1. CIFAR-10 and CIFAR-100 are both widely used datasets in the field of computer vision and machine learning, specifically for image classification tasks. It is evident that our method exhibits greater enhancement over traditional positional encodings on datasets with fewer images per category (CIFAR-100). CIFAR-100 is used for more rigorous testing of models, particularly to evaluate their performance on a larger number of classes and to understand how well they can generalize across a broader range of categories.
Table 1: Comparison with absolute/relative positional encodings on small datasets in image classification benchmarks.
Datasets No-Pos Sinusoidal RoPE LF only RiemannFormer RiemannFormer+LF
CIFAR-10 79.61 86.06 87.91 87.30 89.87 90.82
CIFAR-100 53.37 59.81 62.37 63.01 64.51 66.96
In Table 1, LF is an abbreviation for Locality Focusing mechanism. When employing solely LF without any positional encoding, the module exhibits mediocre performance; however, incorporating LF into positional encoding results in a significant enhancement in model performance. This suggests that LF plays a crucial role in improving performance on these small datasets, whereas positional encoding alone lacks the inherent properties of LF.
4 Conclusion
We assume that the data resides in a curved space, and allow the model to adaptively explore the geometric structure of the space from the data itself, thereby obviating the need for explicit positional information. Our framework can capture both absolute and relative position information and allows the model autonomously balance them during the learning process. Through predefined configurations, we guide the model to capture the geometric structure of curved space while simultaneously reducing the number of parameters that need to be learned. In addition, we introduce a decay factor to attenuate the influence of remote positions. This mechanism facilitates further our modules achieve substantial performance improvements over the baseline.
The framework outlined in this paper has been demonstrated to effectively improve the performance of transformers by conducting experiments on the CIFAR-10 and CIFAR-100 datasets. The current implementation in this paper serves as a preliminary validation of our insights, and further evaluation experiments involving more visual models and large language models will be conducted sequentially.
Appendix
We can also learn a metric matrix of general form, which is not necessarily restricted to the diagonal matrix. For clarity and simplicity, we present here only the $2\times 2$ matrix corresponding to two-dimensional subspace.
Let,
$\mathbf{S}=\left[\begin{matrix}s_{1}&0\\[8.09998pt] 0&s_{2}\end{matrix}\right]$
where $s_{1}>0$ and $s_{2}>0$. Then, the symmetric positive definite matrices are preset as follows,
$\mathbf{M}_{m}$ $=\left[\begin{matrix}\cos{\left(m\theta\right)}&-\sin{\left(m\theta\right)}\\[8.09998pt] \sin{\left(m\theta\right)}&\;\;\;\cos{\left(m\theta\right)}\end{matrix}\right]^{\top}\left[\begin{matrix}s_{1}^{m}&0\\[8.09998pt] 0&s_{2}^{m}\end{matrix}\right]\left[\begin{matrix}\cos{\left(m\theta\right)}&-\sin{\left(m\theta\right)}\\[8.09998pt] \sin{\left(m\theta\right)}&\;\;\;\cos{\left(m\theta\right)}\end{matrix}\right]$
$=\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{m}\cdot\breve{\bf R}(m\theta),$
and,
$\mathbf{M}_{n}$ $=\left[\begin{matrix}\cos{\left(n\theta\right)}&-\sin{\left(n\theta\right)}\\[8.09998pt] \sin{\left(n\theta\right)}&\;\;\;\cos{\left(n\theta\right)}\end{matrix}\right]^{\top}\left[\begin{matrix}s_{1}^{n}&0\\[8.09998pt] 0&s_{2}^{n}\end{matrix}\right]\left[\begin{matrix}\cos{\left(n\theta\right)}&-\sin{\left(n\theta\right)}\\[8.09998pt] \sin{\left(n\theta\right)}&\;\;\;\cos{\left(n\theta\right)}\end{matrix}\right]$
$=\breve{\bf R}(-n\theta)\cdot\mathbf{S}^{n}\cdot\breve{\bf R}(n\theta).$
Then, the linear transformations can be preset as the following matrices,
$\mathbf{T}_{m}$ $=\left[\begin{matrix}\;\;\;\cos{\left(m\theta\right)}&\sin{\left(m\theta\right)}\\[8.09998pt] -\sin{\left(m\theta\right)}&\cos{\left(m\theta\right)}\end{matrix}\right]\left[\begin{matrix}s_{1}^{-\frac{m}{2}}&0\\ 0&s_{2}^{-\frac{m}{2}}\end{matrix}\right]$
$=\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{-\frac{m}{2}},$
and,
$\mathbf{T}_{n}$ $=\left[\begin{matrix}\;\;\;\cos{\left(n\theta\right)}&\sin{\left(n\theta\right)}\\[8.09998pt] -\sin{\left(n\theta\right)}&\cos{\left(n\theta\right)}\end{matrix}\right]\left[\begin{matrix}s_{1}^{-\frac{n}{2}}&0\\ 0&s_{2}^{-\frac{n}{2}}\end{matrix}\right]$
$=\breve{\bf R}(-n\theta)\cdot\mathbf{S}^{-\frac{n}{2}}.$
It is straightforward to verify that:
$\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}=\mathbf{T}_{n}^{\top}\mathbf{M}_{n}\mathbf{T}_{n}.$
and,
$\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$ $=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{m}\cdot\breve{\bf R}(m\theta)\cdot\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{-\frac{m}{2}}\cdot\mathbf{S}^{\frac{n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}$
$=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{\frac{m+n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}$
Then, the form of a standard inner product is,
$\langle s^{\frac{m}{2}}\breve{\bf R}(m\theta)\mathbf{q}_{m},\;s^{\frac{n}{2}}\breve{\bf R}(n\theta)\mathbf{k}_{n}\rangle_{\mathbf{I}}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$