### v2

# RiemannFormer: A Framework for Attention in Curved Spaces

Abstract
This research endeavors to unlock the further potential of transformer-based architectures by providing a geometric interpretation of the self-attention mechanism. In our framework, we reframe attention as interactions on a potential Attention Manifold. Queries and keys are treated as vectors within the manifold’s intrinsic geometry, residing in the tangent bundle, where the geometry of each tangent space is governed by a Riemannian metric. This metric is the fundamental object that defines both the local inner product for similarity computation and the connection that determines parallel transport. In stark contrast, we propose that value vectors are sections of a separate, trivial vector bundle, decoupling the geometric alignment of $Q/K$ from the feature aggregation of $V$. Our primary contribution is a novel, first-principles construction of this parallel transport using Lie group theory, where for 2D data, we derive a positional encoding from the exceptional isomorphism $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$. This allows us to factorize 4D rotations into two commuting isoclinic rotations aligned with the spatial dimensions. This dual-bundle structure provides a principled foundation for our explicit mechanism that highlights a neighborhood by attenuating remote values. Experimental results demonstrate that our modules deliver significant performance improvements over strong baselines, and further evaluations on visual and large language models are underway.
Keywords Transformer $\cdot$ Self-Attention $\cdot$ Curved Space $\cdot$ Tangent Bundle $\cdot$ Fiber Bundle $\cdot$ Parallel Transport
1 Introduction
The transformer architecture, introduced in the seminal paper by Vaswani et al. [Vaswani2017], has revolutionized the field of natural language processing (NLP) and extended its impact to computer vision, audio processing, and other domains. The iterative application of transformer blocks allows tokens to progressively capture higher-level semantic concepts, making the architecture highly effective for tasks requiring contextual understanding.
The transformer block operates on a sequence of tokens, each represented as a vector of fixed dimensionality $d_{\text{model}}$. For a sequence of $L$ tokens, the input to the block is a matrix $X\in\mathbb{R}^{L\times d_{\text{model}}}$, where each row corresponds to a token’s embedding. The transformer block processes this matrix through its core component, Multi-Head Self-Attention (MHSA). The MHSA mechanism is the cornerstone of the transformer, enabling it to aggregate information globally across all tokens in the input sequence. Unlike convolutional neural networks (CNNs), which rely on local receptive fields, or recurrent neural networks (RNNs), which process sequences sequentially, self-attention allows each token to attend to all other tokens simultaneously, capturing long-range dependencies efficiently. Specifically, the attention mechanism computes a score for each pair of tokens by taking the dot product of their query and key vectors, scaled to prevent large values that could destabilize training: $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}$. The dot product $\mathbf{QK}^{\top}\in\mathbb{R}^{L\times L}$ measures the similarity between the query vector of one token and the key vectors of all tokens, producing an attention score matrix. The softmax function normalizes the scores across the key dimension, converting them into weights that sum to $1$ for each query. This ensures that the attention mechanism focuses on the most relevant tokens. The resulting weights are used to compute a weighted sum of the value vectors, producing a new representation for each token that incorporates information from all other tokens. The output of the attention mechanism is a matrix, where each row is a context-aware representation of the corresponding input token. Moreover, the attention mechanism is permutation-equivariant, meaning the output does not depend on the order of the input tokens unless positional information is explicitly added.
Since the self-attention mechanism is permutation-equivariant, it does not inherently capture the order of tokens. To address this, transformers incorporate positional encodings or positional embeddings, which are added to the input token embeddings before processing. Below, we provide a brief overview of positional encodings in transformer. Positional encodings are indispensable in transformer models, embedding sequential order into permutation-equivariant self-attention mechanisms to enable token position modeling in tasks like natural language processing and vision. The seminal sinusoidal positional encodings, introduced by Vaswani et al. [Vaswani2017], remain prevalent due to their deterministic nature and ability to generalize across arbitrary sequence lengths. In contrast, learned positional encodings, as used in BERT [DevlinCLT19], optimize trainable embeddings for task-specific patterns but falter in extrapolating to longer sequences. Relative positional encodings, adopted in Transformer-XL [Dai2019TransformerXLAL], T5 [2020t5], and TUPE [KeHL21], model token distance relationships, enhancing generalization for long-context tasks. Rotary Positional Embeddings (RoPE) efficiently encode relative positions by applying rotation matrices to query and key vectors [Su2024RoPE], with extensions like xPos [sun2023] incorporating dynamic scaling for stable long-sequence training. Recent innovations include ALiBi [alibi], which simplifies computation by introducing attention biases based on token distances. Prior methods like absolute and relative positional encoding are static post-training. Zheng et al. propose Data-Adaptive Positional Encoding (DAPE), which adjusts dynamically based on input context and learned priors [zheng2024]. In addition, Golovneva et al. introduced Contextual Position Encoding (CoPE) [golovneva2024], a novel position encoding method that conditions positions on context by selectively incrementing positions based on model-determined tokens. Although methods like ALiBi and adaptive PEs reduce complexity, they may introduce instability in extremely long sequences, and their generalization across diverse domains remains underexplored.
Since its adaptation to visual tasks with the Vision Transformer (ViT) [DosovitskiyB0WZ21], the transformer has demonstrated exceptional performance across a wide range of applications, from high-level tasks like classification [DeiT2021] and detection [Carion2020] to low-level tasks such as super-resolution and generation. ViTs process images by dividing them into fixed-size patches, which are then treated as tokens similar to words in a sentence. These patches are embedded and fed into a transformer encoder, which leverages self-attention mechanisms to capture global relationships within the image. This approach allows ViTs to learn complex patterns and dependencies across the entire image, leading to impressive performance on various benchmarks. Positional encoding is a crucial component in the application of transformers to visual tasks, enabling the model to retain spatial information about the input image. Unlike natural language processing tasks where the sequence of words carries inherent positional information, images are divided into patches that lack such sequential context. Positional encodings allows the model to account for the order and arrangement of the patches, which is essential for understanding the spatial relationships and structure within the image. The original ViT model uses fixed positional encodings, similar to those used in the original Transformer model for NLP tasks. In addition to learnable and relative positional encodings, alternative schemes have emerged. RoPE excels in language models, particularly in enhancing transformer length extrapolation, yet its potential in computer vision remains largely unexplored. Heo et al. analyzes RoPE’s application to ViT, demonstrating significant improvements in visual tasks like classification, detection and segmentation, with minimal computational overhead [heo2024ropevit]. Chu et al. introduce conditional positional encoding for ViT, dynamically generated based on local input token context [chu2023CPVT]. These approaches potentially provide a more flexible and scalable approach to positional encoding in ViTs.
Despite its strengths, positional encodings require further interpretation and improvement. The mechanistic role of positional information remains unclear, offering a promising direction for future research. To better elucidate the attention mechanism, we endeavor to investigate the deeper implications underlying positional encoding in this paper. In addition, transformers often benefit from being trained on massive datasets. Larger datasets allow transformers to learn more nuanced patterns and relationships in the data, leading to better accuracy and generalization. However, transformers often exhibit suboptimal performance when applied to small datasets. Through experimental analysis, we have found that the introduction of an explicit locality focusing mechanism significantly enhances the performance of transformer in visual tasks.
2 Methodology
Consider a sequence comprising $L$ input tokens, with their embeddings represented as $\{\mathbf{x}_{i}\}_{i=1}^{L}$, where $\mathbf{x}_{i}\in\mathbb{R}^{D}$ denotes the $D$-dimensional embedding vector for the $i$-th token, devoid of positional information. In this paper, we adopt an alternative perspective: we posit the existence of a curved space $\mathcal{M}$ wherein the encoding of a token sequence corresponds to identifying a sequence of points $\{\mathbf{p}_{i}\}$, or a grid of points $\{\mathbf{p}_{ij}\}$ for image data, etc.
2.1 Attention Manifold
We introduce the concept of the Attention Manifold, a novel framework for interpreting Transformer self-attention through non-Euclidean geometry. In this view, token positions are points on a manifold, and their feature embeddings are vectors in the corresponding tangent spaces. The manifold’s geometry is defined by a Riemannian metric, which modulates local interactions and encodes semantic saliency, and a parallel transport operator, which ensures consistent comparisons between vectors at different positions. This geometric formulation provides a first-principles approach to designing attention mechanisms with strong, built-in inductive biases.
Our approach is motivated by a fundamental shift in perspective from an extrinsic to an intrinsic geometric viewpoint. From this standpoint, a manifold is not defined by the absolute coordinates of its points within some ambient space, but rather by the complete set of relationships and transformations that connect them. Consequently, our primary objective is not to parameterize the absolute locations of tokens. Instead, we focus entirely on the crucial task: to directly define a consistent and computationally tractable protocol for the transport of tangent vectors (the feature embeddings) between the tangent spaces at discrete points. In essence, what matters is not where the tokens are, but how their representations can be transformed into one another.
Formally, in differential geometry, the parallel transport of a vector field V{V} along a parameterized curve $c(t)$ is governed by the condition that its covariant derivative along the curve is zero. In local coordinates, this translates to a system of differential equations:
$\frac{dv^{k}}{dt}+\Gamma^{k}_{ij}\frac{dc^{i}}{dt}v^{j}=0,$
where $c(t)$ is the path between points, and $\Gamma^{k}_{ij}$ are the connection coefficients (e.g., Christoffel symbols) that encode the manifold’s curvature.
However, this formal definition presents a fundamental challenge for our discrete and non-parametric setting. We intentionally avoid the complexity and high cost of explicitly defining the manifold’s embedding, the continuous paths $c(t)$ between tokens, or the connection coefficients $\Gamma^{k}_{ij}$. Consequently, this system of equations, while theoretically fundamental, is practically intractable within our framework as we lack the necessary components for its solution. This raises the central question: how can we define a consistent and computationally tractable parallel transport operator without recourse to the underlying continuous geometry? Specifically, the vectors $\mathbf{q}_{i}$ and $\mathbf{k}_{i}$ reside in the tangent spaces $T_{\mathbf{p}_{i}}\mathcal{M}$ at position $\mathbf{p}_{i}$. For instance, $\mathbf{q}_{m}$ and $\mathbf{k}_{m}$ both belong to the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$, whereas $\mathbf{q}_{m}$ and $\mathbf{k}_{n}$ are situated in the tangent spaces at points $\mathbf{p}_{m}$ and $\mathbf{p}_{n}$, respectively. The inner product in the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$ is defined as $\mathbf{x}_{1}^{\top}\mathbf{M}_{m}\mathbf{x}_{2}$, where $\mathbf{M}_{m}$ is a Riemannian metric specifying the inner product structure at $\mathbf{p}_{m}$. Since $\mathbf{k}_{n}$ does not belong to the tangent space at $\mathbf{p}_{m}$, it is not appropriate to directly perform inner products such as $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathbf{k}_{n}$. To address this, a transformation $\mathtt{f}_{n,m}(\mathbf{k}_{n})$ is required to appropriately transport $\mathbf{q}_{n}$ from the tangent space at $\mathbf{p}_{n}$ to that at $\mathbf{p}_{m}$. Only after this transformation can a valid inner product be executed, given by $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathtt{f}_{n,m}(\mathbf{k}_{n})$.
Let us now formalize the attention mechanism within our geometric framework. We consider a discrete sequence of points, $\{\mathbf{p}_{i}\}_{i=1}^{N}$, defined on the Attention Manifold $\mathcal{M}$. The collection of all tangent spaces on the manifold, $\{T_{\mathbf{p}}\mathcal{M}\}_{\mathbf{p}\in\mathcal{M}}$, forms a structure known as the Tangent Bundle, denoted $T\mathcal{M}$. We conceptualize the query and key representations as sections of this tangent bundle. The query and key vector fields, $Q$ and $K$, therefore assign a tangent vector $\mathbf{q}_{i},\mathbf{k}_{i}\in T_{\mathbf{p}_{i}}\mathcal{M}$ to each point $\mathbf{p}_{i}$. These vectors live within the intrinsic, and potentially curved, geometry of the manifold. In stark contrast, the value representations are treated as sections of a Trivial Vector Bundle, $\mathcal{M}\times\mathbb{R}^{D_{v}}$. This construction places the value vectors $\{\mathbf{v}_{i}\}$ in a global, geometry-agnostic Euclidean space. This trivial bundle construction formally decouples the geometric interactions of queries and keys on the manifold from the value representations. A principal advantage of this decoupling is the architectural flexibility it affords: the dimensionality of the value vectors, $D_{v}$, is not required to match the dimensionality of the tangent spaces, $D$, where queries and keys reside. This allows for independent design choices, mirroring the flexibility found in standard Multi-Head Attention architectures. At each point $\mathbf{p}_{i}\in\mathcal{M}$, we sample a specific query and key vector $\mathbf{q}_{i}\in T_{\mathbf{p}_{i}}\mathcal{M}$ and $\mathbf{k}_{i}\in T_{\mathbf{p}_{i}}\mathcal{M}$ from their respective fields, where the tangent space has dimension $D$. The core of the attention mechanism is the interaction between a query vector $\mathbf{q}_{i}$ at point $\mathbf{p}_{i}$ and key vectors from all other points, $\{\mathbf{k}_{j}\}$. To perform a geometrically consistent comparison, each key vector $\mathbf{k}_{j}$ is first transported from its native tangent space $T_{\mathbf{p}_{j}}\mathcal{M}$ to the query’s tangent space $T_{\mathbf{p}_{i}}\mathcal{M}$. The attention scores are then computed via the local inner product.
Finally, these scalar scores are normalized to form attention weights. These weights are then used to compute a weighted sum of the value vectors $\{\mathbf{v}_{j}\}$ in their common Euclidean space $\mathbb{R}^{D_{v}}$. This aggregation, which does not require parallel transport of the values, yields the updated representation at position $\mathbf{p}_{i}$, which itself is a vector in $\mathbb{R}^{D_{v}}$.
2.2 Parallel Transport
In Euclidean space, tangent spaces at various points are isomorphic via translation, enabling straightforward vector transfer. This identification stems from the flat, homogeneous nature of Euclidean geometry, allowing parallel vector movement without additional geometric considerations. On a Riemannian manifold, parallel transport moves vectors along curves, maintaining consistency with intrinsic geometry structure of the manifold during transfer between tangent spaces. To be more specific, the vectors $\mathbf{q}_{m}$ and $\mathbf{k}_{n}$ ($m\neq n$) are in tangent spaces at different positions $\mathbf{p}_{i}$ respectively. Prior to performing the inner product, it is necessary to adapt the vectors to the same space. This can be accomplished, for instance, by applying a parallel transport, denoted as $\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})$, that transports $\mathbf{k}_{n}$ into the tangent space at $\mathbf{q}_{m}$.
Preliminaries: The parallel transport of tangent vectors $\mathbf{v}\in T_{c(\alpha)}\mathcal{M}$ to the tangent space $T_{c(\beta)}\mathcal{M}$ along a curve $c(t):[0,1]\to\mathcal{M}$ is denoted by $\Gamma_{c[\alpha,\beta]}(\mathbf{v})$. Two fundamental properties can be leveraged:
1) $\Gamma_{c[\alpha,\beta]}:\;T_{c(\alpha)}\mathcal{M}\to T_{c(\beta)}\mathcal{M}$ is a linear map.
2) Consider a Riemannian manifold $\mathcal{M}$ equipped with its associated Riemannian connection $\nabla$. Let V{V} be a parallel vector field along a parameterized curve $c(t):[0,1]\to\mathcal{M}$. The vector field V{V} maintains a constant magnitude along the curve $c$.
To summarize, parallel transport provides a way of moving tangent vectors along a curve using the affine connection, and this provides a linear isomorphism between the tangent spaces at the two ends of the curve. However, the isomorphism obtained in this way will in general depend on the choice of the curve. Furthermore, it must be noted that we lack knowledge of the positions of points, curves, and additional structures (such as explicitly defined connection). Consequently, how can we achieve the parallel transport of tangent vectors? In fact, our methodology does not involve the explicit positional embeddings. Instead, we appoint the points and the curves connecting them in the order of tokens and enable the model to learn the linear mappings between tangent spaces at different points. Through these linear mappings, we can perform valid inner product without involving specific positional information.
To formalize the geometric concepts of vector magnitude and the defining property of our parallel transport operator, we can express the squared norm of a tangent vector in both abstract and computational forms. A tangent vector, such as the key vector $\mathbf{k}_{n}$ at position $\mathbf{p}_{n}$, is expressed in a local coordinate basis $\{\partial_{i}\}$ as a linear combination $\sum_{i}k^{i}_{n}\partial_{i}$, where $\{k^{i}_{n}\}$ are its components.
The squared norm of this vector is computed via the inner product defined by the metric tensor $\mathbf{M}_{n}$ at that point. This relationship is articulated by the following equation:
$\underbrace{\langle\sum\nolimits_{i}k^{i}_{n}\partial_{i},\sum\nolimits_{j}k^{j}_{n}\partial_{j}\rangle_{\mathbf{p}_{n}}}_{\text{Abstract Inner Product}}=\underbrace{\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}}_{\text{Computational Form}}=\underbrace{\mathcal{P}_{n\rightarrow m}({\bf k}_{n})^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}({\bf k}_{n})}_{\text{Isometry under Parallel Transport}}.$ (1)
By the property of bilinearity, the inner product expands to $\sum_{i,j}k^{i}_{n}k^{j}_{n}\langle\partial_{i},\partial_{j}\rangle_{\mathbf{p}_{n}}$. The components of the metric tensor $\mathbf{M}_{n}$ are, by definition, the inner products of the basis vectors themselves: $(\mathbf{M}_{n})_{ij}=\langle\partial_{i},\partial_{j}\rangle_{\mathbf{p}_{n}}$. Substituting this definition directly yields the familiar quadratic form $\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}$, where $\mathbf{k}_{n}$ is the column vector of components $\{k^{i}_{n}\}$. This term represents the squared magnitude of the vector $\mathbf{k}_{n}$ according to the local geometry at $\mathbf{p}_{n}$.
(a) Maps between tangent spaces.
First, we introduce a common reference space $\mathbb{R}^{D}$, upon which the linear transformations between tangent spaces at various points are subsequently learned. As shown in Fig. (1(a)), the linear transformation to be learned at point $\mathbf{p}_{i}$ is denoted by $\mathcal{P}_{i}:\mathbb{R}^{D}\to T_{\mathbf{p}_{i}}\mathcal{M}$ and can be represented by a $D\times D$ matrix $\mathbf{T}_{i}$. In this study, we term the transformation $\mathcal{P}_{i}^{-1}$ as tangent mapping or tangent aligning.
Building upon the foregoing analysis, we formally define the parallel transport operator, $\mathcal{P}_{n\rightarrow m}$, which transports a vector $\mathbf{k}_{n}$ from the tangent space at position $n$ to that at position $m$. This is achieved by composing the inverse transformation at the source with the forward transformation at the destination:
$\mathcal{P}_{n\rightarrow m}({\bf k}_{n})=\mathcal{P}_{m}\circ\mathcal{P}^{-1}_{n}({\bf k}_{n})=\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}.$ (2)
This formulation inherently satisfies the property of transitivity. A composition of transport operations from $n$ to $m$ and subsequently from $m$ to $l$ simplifies to a single direct transport from $n$ to $l$. This is demonstrated by the algebraic cancellation of the intermediate operators, which confirms the robustness of our framework and a consistent geometric structure:
$\mathcal{P}_{n\rightarrow m\rightarrow l}({\bf k}_{n})$ $=\mathcal{P}_{l}\circ\mathcal{P}^{-1}_{m}\left(\mathcal{P}_{m}\circ\mathcal{P}^{-1}_{n}({\bf k}_{n})\right)$ (3)
$=\mathbf{T}_{l}\mathbf{T}^{-1}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}$
$=\mathbf{T}_{l}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}$
$=\mathcal{P}_{n\rightarrow l}({\bf k}_{n}).$
Substituting Eq. (2) into Eq. (1), we obtain:
$\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}=\mathbf{k}_{n}^{\top}\left(\mathbf{T}^{-1}_{n}\right)^{\top}\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}.$ (4)
Given the arbitrariness of $\mathbf{k}_{n}$, it consequently follows that:
$\mathbf{M}_{n}=\left(\mathbf{T}^{-1}_{n}\right)^{\top}\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}.$ (5)
Then, simultaneously left-multiplying by matrix $\mathbf{T}_{n}^{\top}$ and right-multiplying by matrix $\mathbf{T}_{n}$, we obtain:
$\mathbf{T}_{n}^{\top}\mathbf{M}_{n}\mathbf{T}_{n}=\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}.$ (6)
The above relationship of these matrices is what we must consider when predefining them. The Riemannian metric $\mathbf{M}_{i}$ is a symmetric positive matrix. To facilitate the model to effectively learn the linear transformations $\mathbf{T}_{i}$, we first decompose them into two components, including scaling and length-preserving in $\mathbb{R}^{D}$: $\mathbf{T}_{i}=\mathbf{S}_{i}\mathbf{R}_{i}$. The introduction of scaling part $\mathbf{S}_{i}$ is motivated by the characteristics of the curved space and the existence of Riemannian metric. The scaling part is represented by a symmetric positive matrix, and its simple form is a diagonal matrix or a scalar matrix $s\mathbf{I}$, where $\mathbf{I}$ is the identity matrix. For instance, we set $\mathbf{S}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}$, then we have $\mathbf{T}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}\mathbf{R}_{i}$. $\mathbf{R}_{i}$ is an orthogonal matrix used for length-preserving, and a rotation is naturally a candidate. A rotation in $n$-dimensional Euclidean space is a linear transformation that preserves lengths of vectors, and it can be represented by an $D\times D$ orthogonal matrix in the special orthogonal group $\mathbf{SO}(D)$.
Given these definitions, it is straightforward to verify that Eq. (6) holds. Furthermore, the metric-induced inner product can be readily converted into the standard inner product as shown in the following formula, thereby simplifying its implementation.
$\langle\mathbf{q}_{m},\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})\rangle_{\mathbf{M}_{m}}$ $=\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$ (7)
$=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{M}_{m}^{-\frac{1}{2}}\mathbf{R}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$
$=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}^{\frac{1}{2}}\mathbf{R}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$
$=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$
To reiterate, the central aim of our work is to empower the model to learn the intrinsic geometry of the Attention Manifold where sequence data reside. This geometry is fully characterized by the fundamental objects: the set of Riemannian metric tensors $\{\mathbf{M}_{m}\}$ that define local inner products and the transformations between tangent spaces. However, a fully general and learnable parameterization for these objects would be computationally intractable and parametrically prohibitive. Therefore, the core of our contribution lies in introducing a principled, parameter-efficient construction for these geometric structures, using predefined configurations derived from Lie group theory.
A Riemannian metric $\mathbf{M}_{m}$ can be represented by a symmetric positive definite matrix, but for the sake of simplifying the problem, we preset it as a symmetric positive definite diagonal matrix:
$\mathbf{M}_{m}=\operatorname{diag}{(s^{(1)}_{m},s^{(2)}_{m},\cdots,s^{(D)}_{m})}$, (8)
where $\langle\partial_{i},\partial_{i}\rangle_{\mathbf{p}_{m}}=s^{(i)}_{m}$, the superscript $(i)$ denotes the component index. To enhance computational efficiency and reduce parametric complexity, we may constrain the metric tensor $\mathbf{M}_{m}$ to be a scalar matrix of the form $\mathbf{M}_{m}=s_{m}\mathbf{I}$, where $s_{m}>0$. This choice is geometrically equivalent to inducing a conformal deformation on the standard Euclidean space $\mathbb{R}^{D}$. In our geometric framework, this matrix $\mathbf{M}_{m}$ is not merely a scaling factor; it is interpreted as the metric tensor that defines the local geometry of the tangent space at position $m$. Its primary function is to modulate the inner product, thereby assigning a notion of semantic importance or saliency to the token at that position. The standard self-attention mechanism implicitly assumes a uniform Euclidean metric ($\mathbf{M}=\mathbf{I}$), treating all positions as equally important a priori. By introducing a position-dependent metric $\mathbf{M}_{m}$, we allow the model to differentiate the intrinsic significance of tokens. A larger scalar $s_{m}$ effectively amplifies the length of the query vector $\mathbf{q}_{m}$ and, consequently, the magnitude of its computed attention scores to all other keys. Thus, $s_{m}$ acts as a control knob for the token’s influence: a higher value signifies greater semantic importance, causing its query to have a stronger impact on the overall attention distribution.
While the general formulation $\mathbf{M}_{m}=s_{m}\mathbf{I}$ allows for arbitrary, position-specific importance, it does not explicitly encode a structural bias related to sequential order. To introduce such a prior, we can parameterize the scaling factor as an exponential function of the position index $m$. This is elegantly achieved by setting $\mathbf{M}_{m}=s^{m}\mathbf{I}$, where $s>0$ is a single, globally shared, learnable parameter. This highly parsimonious choice allows the model to learn a global decay rate ($s<1$) or a growth rate ($s>1$) of semantic importance, thereby systematically weighting tokens based on their absolute position in the sequence. We refer to this scheme as Conformal Exponential Scaling (CES). To impose the constraint $s<1$, we may proceed by defining it as $s=\frac{\exp(w)}{\exp(w)+\alpha}$, where $w$ is a learnable parameter and $\alpha$ denotes a hyperparameter. Alternatively, one might opt for $s=\exp(w)$, which permits $s>1$. For long sequences, we can decelerate the rate of exponential growth, by setting the parameter $s^{\log_{\beta}(m)}$. In addition, the learnable metric matrix of general form is given in the Appendix section.
2.3 A Principled Decomposition of High-Dimensional Rotations
A rotation in $\mathbb{R}^{D}$ can be represented by a $D\times D$ matrix in the special orthogonal group $\mathbf{SO}(D)$. It is well known that an element of $\mathbf{SO}(D)$ can be generated from the exponential of a skew-symmetric matrix $X$, denoted as $\exp(X)$. Specifically, the orthogonal Lie algebra $\mathfrak{so}(D=2k)$:
$\{X\mid X^{\top}+X=0\}$
yields $D$-dimensional special orthogonal matrices via the exponential map: $\exp(X)\in\mathbf{SO}(D)$.
For position $i$, we predefine the linear transformation as,
$\mathbf{T}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}\exp(iX).$ (9)
Subsequently, we obtain,
$\mathbf{T}_{m}\mathbf{T}_{n}^{-1}$ $=\mathbf{M}_{m}^{-\frac{1}{2}}\exp(m{X})\exp(-n{X})\mathbf{M}_{n}^{\frac{1}{2}}$ (10)
$=\mathbf{M}_{m}^{-\frac{1}{2}}\exp((m-n){X})\mathbf{M}_{n}^{\frac{1}{2}},$
and
$\langle\mathbf{q}_{m},\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})\rangle_{\mathbf{M}_{m}}=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$ (11)
It is worth noting that antisymmetric matrices possess D(D1)/2{D(D-1)}/{2} independent elements, thereby introducing a substantial number of learnable parameters. To reduce the number of learnable parameters and mitigate algorithmic complexity, we may alternatively learn subgroups of the orthogonal matrix group $\mathbf{SO}(D)$, such as Maximal Torus $(S^{1})^{k}$, or subspaces of the orthogonal Lie algebra, such as Cartan subalgebras.
We uniformly designate the encoding scheme herein as Parallel Transport of Embeddings (PaTE). When emphasizing a specific subspace, such as SO(4), it is denoted as PaTE-SO(4).
2.3.1 Decomposition into Commuting $\mathbf{SO}(2)$ Subspaces
We can construct a basis for the Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$. This basis consists of $k$ matrices $\{H_{1},H_{2},\dots,H_{k}\}$:
$H_{1}=\begin{bmatrix}J&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad H_{2}=\begin{bmatrix}0&0&\cdots&0\\ 0&J&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad\dots\quad H_{k}=\begin{bmatrix}0&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&J\end{bmatrix}$,
where each $H_{i}$ is a $2k\times 2k$ block matrix, with all blocks being zero matrices except for the $i$-th diagonal block, which is the $2\times 2$ antisymmetric matrix $J$. The matrix $J$ is the generator of infinitesimal rotations in two dimensions:
$J=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$.
The Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$ is the linear span of this basis $\{H_{1},H_{2},\dots,H_{k}\}$. Any element $H$ within it can be expressed as:
$H=\sum_{i=1}^{k}\theta_{i}H_{i}=\operatorname{diag}{(\theta_{1}J,\theta_{2}J,\dots,\theta_{k}J)},$
where $\theta_{1},\theta_{2},\dots,\theta_{k}$ are arbitrary real numbers, and $\operatorname{diag}({B}_{1},{B}_{2},\dots,{B}_{n})$ denotes the block diagonal matrix with each ${B}_{i}$ serving as a diagonal block. The matrix $H$ represents simultaneous infinitesimal rotations on $k$ orthogonal planes, with the "angular velocity" on each plane being $\theta_{i}$.
The exponential of $H$ is
$\exp(H)=\exp(\sum_{i=1}^{k}\theta_{i}H_{i})=\operatorname{diag}{(\exp(\theta_{1}J),\exp(\theta_{2}J),\dots,\exp(\theta_{k}J))},$
where a diagonal block $\exp(\theta_{i}{J})$ is a $2\times 2$ rotation matrix,
$\exp(\theta_{i}{J})=\cos(\theta_{i}){I}+\sin(\theta_{i}){J}=\begin{bmatrix}\cos(\theta_{i})&\!\!-\sin(\theta_{i})\\ \sin(\theta_{i})&\;\;\cos(\theta_{i})\end{bmatrix}$.
The interaction between two positions $m$ and $n$ can be represented as
$\exp\left(m\theta{J}\right)\exp\left(n\theta{J}\right)^{-1}$ $=\exp\left((m-n)\theta{J}\right)$
$=\cos\left((m-n)\theta\right){I}+\sin\left((m-n)\theta\right){J}$
$=\begin{bmatrix}\cos((m-n)\theta)&\!\!-\sin((m-n)\theta)\\ \sin((m-n)\theta)&\;\;\cos((m-n)\theta)\end{bmatrix}$.
In fact, for any real antisymmetric matrix $X$, there exists an orthogonal matrix $Q$ such that $X=QHQ^{T}$, where $H$ is an element in the Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$, and thus $\exp(X)=Q\exp(H)Q^{T}$. Here, the rotation in the 2D subspace corresponds to setting $Q=I$, which simplifies the scenario but also limits its capacity for interactions across spatial dimensions. It is also worth noting that, under the aforementioned configuration, the proposed approach reduces to a conventional RoPE when setting the metric as $\mathbf{M}_{i}=\mathbf{I}$, thereby indicating that RoPE can be viewed as a special case of our framework.
To enhance interactions across more dimensions, one may consider rotational transformations within higher-dimensional subspaces.
2.3.2 $\mathbf{SO}(3)$ Subspaces
We can consider rotational transformations within 3-dimensional subspaces. We assume here that $D$ is a multiple of 3, i.e., $D=3k$.
First, we define the $3\times 3$ unit antisymmetric matrix $K$ corresponding to a unit rotation axis:
$K=\begin{bmatrix}0&-c&b\\ c&0&-a\\ -b&a&0\end{bmatrix}$,
where $a^{2}+b^{2}+c^{2}=1$, i.e., $\mathbf{v}=[a,b,c]^{T}$ is a unit vector representing the rotation axis. For a real number $\theta$, the exponential of the antisymmetric matrix $\theta K$ is:
$\exp(\theta K)=I+\sin(\theta)K+(1-\cos(\theta))K^{2}.$
This is the classic Rodrigues’ rotation formula, which gives a closed-form exponential map from the Lie algebra $\mathfrak{so}(3)$ to the Lie group $\mathbf{SO}(3)$, transforming an antisymmetric matrix encoding the rotation axis and angle into its corresponding rotation matrix.
Incorporating positional information:
$\exp(m\theta K)=I+\sin(m\theta)K+(1-\cos(m\theta))K^{2}.$
Likewise, the interaction between two positions $m$ and $n$ can be represented as
$\exp\left(m\theta K\right)\exp\left(n\theta K\right)^{-1}$ $=\exp\left((m-n)\theta K\right)$
$=I+\sin\left((m-n)\theta\right)K+(1-\cos\left((m-n)\theta)\right)K^{2}.$
2.3.3 The Isoclinic Decomposition of $\mathbf{SO}(4)$
It is well-established that rotations in the special orthogonal group $\mathbf{SO}(4)$ are generated via the exponential map from its Lie algebra, $\mathfrak{so}(4)$. Elements $X\in\mathfrak{so}(4)$ are $4\times 4$ real skew-symmetric matrices of the form:
$X=\begin{bmatrix}0&x_{1}&x_{2}&x_{3}\\ -x_{1}&0&x_{4}&x_{5}\\ -x_{2}&-x_{4}&0&x_{6}\\ -x_{3}&-x_{5}&-x_{6}&0\end{bmatrix}$.
However, for an arbitrary $X$, the direct computation of its matrix exponential, $\exp(X)$, is often complex and lacks a concise closed-form expression analogous to that in 2D, such as $\exp(\theta J)=\cos(\theta)I+\sin(\theta)J$ where $J^{2}=-I$. To devise a more efficient and insightful computational framework, it is imperative to investigate the intrinsic algebraic structure of $\mathfrak{so}(4)$.
The key insight lies in an exceptional property of the $\mathfrak{so}(4)$ Lie algebra: it is the unique semi-simple Lie algebra that decomposes into a direct sum of two ideals. Specifically, $\mathfrak{so}(4)$ is isomorphic to the direct sum of two $\mathfrak{su}(2)$ Lie algebras:
$\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2).$
Given the isomorphism between $\mathfrak{su}(2)$ and the Lie algebra of 3D rotations, $\mathfrak{so}(3)\cong\mathfrak{su}(2)$, the above relation leads to a more intuitive isomorphism over the real field: $\mathfrak{so}(4)\cong\mathfrak{so}(3)\oplus\mathfrak{so}(3)$. This profound structural property implies that the algebra of 4D rotations can be perfectly deconstructed into a combination of two independent, familiar $\mathfrak{so}(3)$ algebras. This foundational decomposition provides the basis for simplifying the exponential map by decomposing $\mathfrak{so}(4)$ into a direct sum of two subalgebras.
To realize this decomposition, we define two special sets of basis vectors in the six-dimensional vector space of $\mathfrak{so}(4)$, denoted $\{A_{1},A_{2},A_{3}\}$ and $\{B_{1},B_{2},B_{3}\}$. These bases respectively span two three-dimensional subalgebras, $\mathfrak{h}_{1}=\text{span}\{A_{1},A_{2},A_{3}\}$ and $\mathfrak{h}_{2}=\text{span}\{B_{1},B_{2},B_{3}\}$. Since their union forms a complete basis for $\mathfrak{so}(4)$, any skew-symmetric matrix $X$ can be uniquely decomposed as $X=A+B$, where $A=\sum a_{i}A_{i}\in\mathfrak{h}_{1}$ and $B=\sum b_{j}B_{j}\in\mathfrak{h}_{2}$. Thus, $\mathfrak{so}(4)$ is the direct sum of its subalgebras $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$, denoted $\mathfrak{so}(4)=\mathfrak{h}_{1}\oplus\mathfrak{h}_{2}$.
The principal advantage of this decomposition is the simplification of the exponential map. As any element $A\in\mathfrak{h}_{1}$ commutes with any element $B\in\mathfrak{h}_{2}$, the exponential of their sum factors into a product of exponentials:
$\exp(X)=\exp(A+B)=\exp(A)\exp(B).$
This reduces the computation for a complex 4D rotation to that of two independent, quasi-3D rotations, each admitting an efficient, closed-form solution.
The standard bases for implementing the decomposition $\mathfrak{so}(4)=\mathfrak{h}_{1}\oplus\mathfrak{h}_{2}$ are defined as follows. The basis $\{A_{1},A_{2},A_{3}\}$ for the subalgebra $\mathfrak{h}_{1}$ is:
$A_{1}=\begin{bmatrix}0&0&0&-1\\ 0&0&-1&0\\ 0&1&0&0\\ 1&0&0&0\end{bmatrix},\;A_{2}=\begin{bmatrix}0&0&1&0\\ 0&0&0&-1\\ -1&0&0&0\\ 0&1&0&0\end{bmatrix},\;A_{3}=\begin{bmatrix}0&-1&0&0\\ 1&0&0&0\\ 0&0&0&-1\\ 0&0&1&0\end{bmatrix}$.
Any element in $\mathfrak{h}_{1}$ is a linear combination of these bases. Consider a normalized generator $J_{A}$ defined by a unit vector $(a_{1},a_{2},a_{3})$ such that $a_{1}^{2}+a_{2}^{2}+a_{3}^{2}=1$:
$J_{A}=\sum_{i=1}^{3}a_{i}A_{i}=\begin{bmatrix}0&-a_{3}&a_{2}&-a_{1}\\ a_{3}&0&-a_{1}&-a_{2}\\ -a_{2}&a_{1}&0&-a_{3}\\ a_{1}&a_{2}&a_{3}&0\end{bmatrix}$.
By construction, $J_{A}$ satisfies $J_{A}^{2}=-I$, where $I$ is the identity matrix. Its algebraic behavior is therefore analogous to the imaginary unit $i$. Consequently, its matrix exponential, derived from the Taylor series expansion, assumes a form akin to Euler’s formula:
$\exp(\theta_{1}J_{A})=\cos(\theta_{1}){I}+\sin(\theta_{1})J_{A}.$
Expanding this expression yields an explicit rotation matrix in $\mathbf{SO}(4)$:
$\exp(\theta_{1}J_{A})=\begin{bmatrix}\cos(\theta_{1})&-a_{3}\sin(\theta_{1})&a_{2}\sin(\theta_{1})&-a_{1}\sin(\theta_{1})\\ a_{3}\sin(\theta_{1})&\cos(\theta_{1})&-a_{1}\sin(\theta_{1})&-a_{2}\sin(\theta_{1})\\ -a_{2}\sin(\theta_{1})&a_{1}\sin(\theta_{1})&\cos(\theta_{1})&-a_{3}\sin(\theta_{1})\\ a_{1}\sin(\theta_{1})&a_{2}\sin(\theta_{1})&a_{3}\sin(\theta_{1})&\cos(\theta_{1})\end{bmatrix}$,
where $\theta_{1}$ represents the angle of rotation about an axis defined by the vector $(a_{1},a_{2},a_{3})$.
In a completely analogous manner, we define the second basis $\{B_{1},B_{2},B_{3}\}$, which spans the commuting subalgebra $\mathfrak{h}_{2}$:
$B_{1}=\begin{bmatrix}0&0&0&1\\ 0&0&-1&0\\ 0&1&0&0\\ -1&0&0&0\end{bmatrix},\;B_{2}=\begin{bmatrix}0&0&1&0\\ 0&0&0&1\\ -1&0&0&0\\ 0&-1&0&0\end{bmatrix},\;B_{3}=\begin{bmatrix}0&-1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&-1&0\end{bmatrix}$.
We construct a normalized generator $J_{B}$ from a unit vector $(b_{1},b_{2},b_{3})$:
$J_{B}=\sum_{i=1}^{3}b_{i}B_{i}=\begin{bmatrix}0&-b_{3}&b_{2}&b_{1}\\ b_{3}&0&-b_{1}&b_{2}\\ -b_{2}&b_{1}&0&b_{3}\\ -b_{1}&-b_{2}&-b_{3}&0\end{bmatrix}$,
where $b_{1}^{2}+b_{2}^{2}+b_{3}^{2}=1$. The basis $\{B_{i}\}$ shares the same algebraic properties as $\{A_{i}\}$, leading to $J_{B}^{2}=-I$ and the identical form for its exponential map:
$\exp(\theta_{2}J_{B})=\cos(\theta_{2}){I}+\sin(\theta_{2})J_{B},$
where $\theta_{2}$ is another independent angle of rotation.
Rotations generated by $A\in\mathfrak{h}_{1}$, such as $\exp(A)$, are termed left-isoclinic rotations, while those generated by $B\in\mathfrak{h}_{2}$ are right-isoclinic rotations. Any rotation in 4D can be uniquely decomposed into the product of a left-isoclinic and a right-isoclinic rotation, which commute. This gives rise to the exponential property $\exp(A+B)=\exp(A)\exp(B)$ for $A\in\mathfrak{h}_{1},\,B\in\mathfrak{h}_{2}$.
We introduce rotation angles $\theta_{A}$ and $\theta_{B}$ for the two types of rotations. The rotation matrices corresponding to positions $m$ and $n$ are then respectively defined as:
$R(\theta_{A},\theta_{B},m)=\exp\left(m\theta_{A}J_{A}\right)\exp\left(m\theta_{B}J_{B}\right)\quad\text{and}\quad R(\theta_{A},\theta_{B},n)=\exp\left(n\theta_{A}J_{A}\right)\exp\left(n\theta_{B}J_{B}\right).$
The interaction between two positions $m$ and $n$ is captured by the relative transformation:
$R(\theta_{A},\theta_{B},m)R(\theta_{A},\theta_{B},n)^{-1}.$
For a single generator, e.g., $J_{A}$, the relative transformation simplifies to:
$\exp\left(m\theta J_{A}\right)\exp\left(n\theta J_{A}\right)^{-1}=\exp((m-n)\theta J_{A})=\cos((m-n)\theta){I}+\sin((m-n)\theta)J_{A}.$
An identical relation holds for the matrix $J_{B}$. Since the Lie bracket $J_{A},J_{B}]=0$, their exponentials commute, allowing us to write:
$R(\theta_{A},\theta_{B},m)R(\theta_{A},\theta_{B},n)^{-1}=\exp\left((m-n)\theta_{A}J_{A}\right)\exp\left((m-n)\theta_{B}J_{B}\right).$
When the feature dimension $D$ is a multiple of $4k$, we can apply this method to perform rotations in each decoupled 4D subspace. For a position $i$, we define the transformation matrix as:
$\mathbf{T}_{i}=s_{i}^{-1}R(\theta_{A},\theta_{B},i)=s_{i}^{-1}\exp\left(i\theta_{A}J_{A}\right)\exp\left(i\theta_{B}J_{B}\right).$
The interaction between two positions $m$ and $n$ corresponds to the following matrix:
$\mathbf{T}_{m}{\mathbf{T}_{n}}^{\!\!-1}$ $=s_{m}^{-1}R(\theta_{A},\theta_{B},m)s_{n}R(\theta_{A},\theta_{B},n)^{-1}$
$={\frac{s_{n}}{s_{m}}}\exp\left((m-n)\theta_{A}J_{A}\right)\exp\left((m-n)\theta_{B}J_{B}\right).$
2D PaTE-SO(4)
Notably, this approach is applicable not only to 1D token sequences, as in language models, but also to 2D sequences of image patches. We can associate $J_{A}\in\mathfrak{h}_{1}$ with the $x$-axis and $J_{B}\in\mathfrak{h}_{2}$ with the $y$-axis:
$R(\theta_{A},\theta_{B},(x_{1},y_{1}))$ $=\exp\left(x_{1}\theta_{A}J_{A}\right)\exp\left(y_{1}\theta_{B}J_{B}\right),$
$R(\theta_{A},\theta_{B},(x_{2},y_{2}))$ $=\exp\left(x_{2}\theta_{A}J_{A}\right)\exp\left(y_{2}\theta_{B}J_{B}\right).$
For two patches located at $(x_{1},y_{1})$ and $(x_{2},y_{2})$, the relative transformation is:
$\mathbf{T}_{(x_{1},y_{1})}{\mathbf{T}_{(x_{2},y_{2})}}^{\!-1}$ $=s_{x_{1}}^{-1}s_{y_{1}}^{-1}R(\theta_{A},\theta_{B},(x_{1},y_{1}))s_{x_{2}}s_{y_{2}}R(\theta_{A},\theta_{B},(x_{2},y_{2}))^{-1}$
$={\frac{s_{x_{2}}s_{y_{2}}}{s_{x_{1}}s_{y_{1}}}}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right).$
If we define $\mathbf{M}_{(x_{1},y_{1})}=s^{2}_{x_{1}}s^{2}_{y_{1}}\mathbf{I}$, then the interaction term can be expressed as:
$\mathbf{q}_{\mathbf{m}}^{\top}\mathbf{M}_{\mathbf{m}}\mathbf{T}_{\mathbf{m}}{\mathbf{T}_{\mathbf{n}}}^{\!-1}\mathbf{k}_{\mathbf{n}}$ $=s_{x_{1}}s_{y_{1}}s_{x_{2}}s_{y_{2}}\mathbf{q}_{m}^{\top}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right)\mathbf{k}_{\mathbf{n}}$
$=\langle\mathbf{T}_{\mathbf{m}}^{-1}\mathbf{q}_{\mathbf{m}},\;{\mathbf{T}_{\mathbf{n}}}^{\!-1}\mathbf{k}_{\mathbf{n}}\rangle,$
where $\mathbf{m}$ and $\mathbf{n}$ denote 2D positions $(x_{1},y_{1})$ and $(x_{2},y_{2})$ for brevity in the expression. The final equation provides a concrete instantiation of our geometric framework for 2D data. The relative transformation is a composition of two orthogonal transformations operating in distinct, commuting subspaces, perfectly encoding the displacement along each spatial axis. This formulation is inherently isotropic, treating the $x$ and $y$ dimensions symmetrically, yet allows for learned anisotropic modulation through the scaling factors $s_{x}$ and $s_{y}$. Ultimately, this method yields a holistic representation of 2D relative positions, standing in contrast to methods that treat spatial dimensions as separate or unstructured features. The unique algebraic structure of $\mathfrak{so}(4)$, specifically its decomposition into two commuting subalgebras, finds a natural correspondence with the geometry of the 2D plane, resulting in a computationally elegant and theoretically robust representation of spatial relationships.
2.4 Inducing Structure on the Trivial Value Bundle
The formalization of value vectors as sections of a trivial bundle, $\mathcal{M}\ \times\mathbb{R}^{D_{v}}$, is not merely a descriptive choice; it is a generative one. It provides a principled foundation for introducing more sophisticated, structured operations beyond simple weighted averaging. We explore three such extensions that leverage the fiber bundle structure to enhance the expressive power of the attention mechanism.
Dynamic Fiber Gating
The standard attention mechanism employs a scalar weight, $a_{m,i}$, which uniformly scales all channels of a value vector $\mathbf{v}_{i}$. This raises a natural question: should the query $\mathbf{q}_{m}$ only determine how much information to retrieve from position $i$, or should it also influence what kind of information is retrieved?
To address this, we introduce a dynamic fiber gating mechanism. We define a learnable gating function, $g$, that is conditioned on the query vector and acts directly on the value fiber $\mathbb{R}^{D_{v}}$. This function maps the query’s tangent space to a gating vector in the value space:
$g:T_{\mathbf{p}_{m}}\mathcal{M}\rightarrow\mathbb{R}^{D_{v}}.$
The aggregation of values is then modulated by this gate, allowing for content-addressable channel selection:
$\mathbf{o}_{m}=\sum_{i}a_{m,i}\left(g(\mathbf{q}_{m})\odot\mathbf{v}_{i}\right),$
where $\odot$ denotes the element-wise product. This mechanism empowers the model to dynamically select and emphasize the most relevant feature channels from a value vector, conditioned on the specific query.
Local Symmetry via Group Action
The trivial bundle construction assumes that value vectors can be directly aggregated. However, for data with inherent geometric properties, it may be beneficial to apply a position-dependent transformation to "align" features before aggregation.
This can be formalized by introducing a group action on the value fiber. Let $G$ be a Lie group acting on $\mathbb{R}^{D_{v}}$, such as the special orthogonal group $\mathbf{SO}(D_{v})$. We can then learn a mapping, $h$, from the base manifold to the group itself:
$h:\mathcal{M}\rightarrow G.$
The value aggregation is then modified to incorporate these local symmetry transformations:
$\mathbf{o}_{m}=\sum_{i}a_{m,i}\left(h(\mathbf{p}_{i})\cdot\mathbf{v}_{i}\right),$
where $h(\mathbf{p}_{i})\cdot\mathbf{v}_{i}$ represents the action of the group element $h(\mathbf{p}_{i})$ (e.g., a rotation matrix) on the vector $\mathbf{v}_{i}$. This introduces a powerful inductive bias, often referred to as a local gauge symmetry, allowing the model to learn position-aware feature adaptations.
Connections on the Value Bundle
Our initial assumption of a trivial bundle implies that the "rules" for aggregating values are globally uniform and position-agnostic. We can challenge this assumption by introducing a new, learnable connection on the value bundle.
A connection defines a notion of parallel transport specific to the bundle it is on. Let us denote this new value connection as $\mathcal{C}$, with its corresponding parallel transport operator $\mathcal{C}_{i\rightarrow m}$. Crucially, this connection is independent of the Riemannian connection on the tangent bundle used for $Q/K$ interactions. The final aggregation step then becomes:
$\mathbf{o}_{m}=\sum_{i}a_{m,i}\,\mathcal{C}_{i\rightarrow m}(\mathbf{v}_{i}).$
This establishes a dual-geometry framework:
1) A Riemannian geometry on the tangent bundle governs the $Q/K$ interactions, determining where to attend.
2) A gauge-like geometry, defined by the learned value connection $\mathcal{C}$, governs the $V$ aggregation, determining how to integrate the attended information.
This provides maximal flexibility, allowing the model to learn, for instance, that information from distant tokens should be transformed (e.g., blurred or down-weighted) differently than information from nearby tokens during the final value aggregation.
A Concrete Realization: The Locality Focusing Mechanism
While the concepts of dynamic gating, group actions, and fully learned connections represent powerful avenues for future research, we now introduce a practical and highly effective mechanism that realizes the core principle of a structured value bundle with minimal overhead. This approach can be understood as a specific instantiation of a predefined connection on the value bundle. Instead of a complex learnable operator, this connection is simplified to a position-dependent scalar modulation, designed explicitly to enforce a local inductive bias. We term this mechanism Locality Focusing.
Unlike CNNs, which are endowed with locality and translational equivariance, vanilla transformers are data-hungry due to their minimal inductive biases. This flexibility makes them powerful on large, diverse datasets but can be a limitation on limited data, where domain-specific priors are crucial for generalization. Our Locality Focusing mechanism directly addresses this by introducing an explicit neighborhood bias, attenuating the influence of remote values during aggregation.
Consistent with the trivial bundle structure that places value vectors in a global Euclidean space $\mathbb{R}^{D}$, this modulation is performed directly on the values prior to their final weighting. While attention scores implicitly reflect semantic relevance, Locality Focusing enforces a strong spatial bias. Specifically, for a given query position $m$, we modulate each value vector $\mathbf{v}_{i}$ by a position-relative decay factor, $\lambda_{m,i}$. This factor is designed to be a monotonic function of the distance between positions $m$ and $i$. The final output vector for position $m$ is then computed as the weighted sum of these modulated values, where $s_{m,i}$ are the softmax-normalized attention scores:
$\mathbf{o}_{m}=\sum_{i}s_{m,i}\left(\lambda_{m,i}\mathbf{v}_{i}\right).$
This approach effectively disentangles the "what" (semantic relevance from attention) from the "where" (spatial locality from the $\lambda$ factors), providing a more robust and controllable inductive bias. In matrix form, the output for a single query $\mathbf{q}_{m}$ is:
$\mathbf{o}_{m}=\text{Attention}(\mathbf{q}_{m},\mathbf{K},\mathbf{\Lambda}_{m},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{q}_{m}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)(\mathbf{\Lambda}_{m}\mathbf{V})=\mathbf{s}_{m}(\mathbf{\Lambda}_{m}\mathbf{V}),$ (12)
where $\mathbf{s}_{m}\in\mathbb{R}^{1\times L}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times D}$, and $\mathbf{\Lambda}_{m}\in\mathbb{R}^{L\times L}$ is a diagonal matrix containing the factors $\{\lambda_{m,i}\}_{i=1}^{L}$. This expands to:
$\mathbf{o}_{m}$ $=\begin{bmatrix}s_{m,1}&s_{m,2}&\cdots&s_{m,L}\end{bmatrix}\begin{bmatrix}\lambda_{m,1}&&&\\ &\lambda_{m,2}&&\\ &&\ddots&\\ &&&\lambda_{m,L}\end{bmatrix}\begin{bmatrix}\mathbf{v}_{1}\\[3.6pt] \mathbf{v}_{2}\\ \vdots\\ \mathbf{v}_{L}\end{bmatrix}$ (13)
$=\begin{bmatrix}s_{m,1}&s_{m,2}&\cdots&s_{m,L}\end{bmatrix}\begin{bmatrix}\lambda_{m,1}\mathbf{v}_{1}\\[3.6pt] \lambda_{m,2}\mathbf{v}_{2}\\ \vdots\\ \lambda_{m,L}\mathbf{v}_{L}\end{bmatrix}$.
This mechanism, while presented here for 2D image data, is equally applicable to n-dimensional sequences. Its formulation is strongly reminiscent of the bilateral filter from image processing, which smooths images while preserving edges:
$\mathbf{o}_{m}=\sum_{n=1}^{L}\underbrace{G_{s}(\mathbf{q}_{m},\mathbf{k}_{n})}_{\text{Intensity Similarity}}\underbrace{G_{\sigma}(\mathbf{d}_{m},\mathbf{d}_{n})}_{\text{Spatial Proximity}}\mathbf{v}_{n},$ (14)
Here, the attention term $G_{s}(\cdot,\cdot)$ plays the role of the intensity/range kernel, measuring similarity in feature space, while our attenuation factor $G_{\sigma}(\cdot,\cdot)$ acts as the spatial/domain kernel. A common choice for $G_{\sigma}$ is a Gaussian function of the distance between positions $\mathbf{d}_{m}$ and $\mathbf{d}_{n}$, $G_{\sigma}(\mathbf{x},\mathbf{y})=e^{-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}_{A}}{2\sigma^{2}}}$, where $\mathbf{A}$ and $\sigma$ can be learnable.
For efficient implementation across all tokens, the attenuation factors for all query positions can be collected into a single $L\times L$ matrix $\mathbf{\Omega}$, where $\mathbf{\Omega}_{m,i}=\lambda_{m,i}$. The full attention output is then computed as:
$\mathbf{O}=(\mathbf{S}\odot\mathbf{\Omega})\mathbf{V}=\left(\begin{bmatrix}s_{1,1}&s_{1,2}&\cdots&s_{1,L}\\ s_{2,1}&s_{2,2}&\cdots&s_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ s_{L,1}&s_{L,2}&\cdots&s_{L,L}\end{bmatrix}\odot\begin{bmatrix}\lambda_{1,1}&\lambda_{1,2}&\cdots&\lambda_{1,L}\\ \lambda_{2,1}&\lambda_{2,2}&\cdots&\lambda_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{L,1}&\lambda_{L,2}&\cdots&\lambda_{L,L}\end{bmatrix}\right)\cdot\mathbf{V}.$ (15)
This simple mechanism significantly improves the capacity of transformers to learn the inductive bias of locality, especially on smaller datasets. As our experiments reveal, incorporating such an explicit local focusing mechanism substantially improves the performance of transformers in visual tasks.
3 Experiments
This study primarily focuses on the visual domain and, consequently, evaluates the performance of RiemnnFormer on image classification tasks for small datasets.
In the context of image data, there are two axes to consider. Although these axes can be addressed jointly [heo2024ropevit] or the image patches are directly flatten into a sequence, we are interested in exploring the effects of processing them separately along the $x$ and $y$ axes. This strategy also facilitates a comparative studies with previous work.
The proposed modules undergo training and evaluation on the CIFAR-10 and CIFAR-100 datasets, employing a lightweight Vision Transformer (ViT) architecture. Specifically, these modules leverage the compact yet efficient structure of the ViT model to process the image classification tasks inherent to these benchmark datasets.
All experiments were implemented on an NVIDIA RTX 4090 GPU with PyTorch 2.7. The training procedure is conducted over 200 epochs with a batch size of 128, utilizing the AdamW optimizer configured with its default hyperparameters.
In predefined transformations, rotation operations in subspaces involve the configuration of angles. An angle schedule is a function that sends an integer to an angle involved in the linear transformation. The choice of the angle schedule can also impact the performance of the tangent mapping. In our experiments, we observed that the following straightforward schedule yields favorable results: $f(t):t\mapsto t\theta,\;t\in\{1,2,\dots\}$, where $\theta$ is a learnable parameter or a predefined angle. Exploring more effective angle schedules is also a direction worth investigating. In the current experiment of this work, we use 1D PaTE-SO(2) with the Conformal Exponential Scaling scheme ($\mathbf{M}_{i}=s^{i}\mathbf{I}$), and we set the hyperparameter $\alpha=0.1$ in $s=\frac{\exp(w)}{\exp(w)+\alpha}$ in these experiments.
The current work is primarily dedicated to further exploring the potential of transformer-based models in visual tasks, and thus, it mainly compares with ViT architectures. The architecture of our ultra-lightweight vision transformer comprises only $4$ attention heads, $6$ transformer layers, and decomposes images into $64$ patches, each of $4\times 4$ pixels. The results from the experiments are compiled in Table 1. CIFAR-10 and CIFAR-100 are both widely used datasets in the field of computer vision and machine learning, specifically for image classification tasks. It is evident that our method exhibits greater enhancement over traditional positional encodings on datasets with fewer images per category (CIFAR-100). CIFAR-100 is used for more rigorous testing of models, particularly to evaluate their performance on a larger number of classes and to understand how well they can generalize across a broader range of categories.
Table 1: Comparison with absolute/relative positional encodings on small datasets in image classification benchmarks.
Datasets No-Pos Sinusoidal RoPE LF only RiemannFormer RiemannFormer+LF
CIFAR-10 79.61 86.06 87.91 87.30 89.87 90.82
CIFAR-100 53.37 59.81 62.37 63.01 64.51 66.96
In Table 1, LF is an abbreviation for Locality Focusing mechanism. When employing solely LF without any positional encoding, the module exhibits mediocre performance; however, incorporating LF into positional encoding results in a significant enhancement in model performance. This suggests that LF plays a crucial role in improving performance on these small datasets, whereas positional encoding alone lacks the inherent properties of LF.
4 Conclusion
We assume that the data resides in a curved space, and allow the model to adaptively explore the geometric structure of the space from the data itself, thereby obviating the need for explicit positional information. Our framework can capture both absolute and relative position information and allows the model autonomously balance them during the learning process. Through predefined configurations, we guide the model to capture the geometric structure of curved space while simultaneously reducing the number of parameters that need to be learned. In addition, we introduce a decay factor to attenuate the influence of remote positions. This mechanism facilitates further our modules achieve substantial performance improvements over the baseline.
The framework outlined in this paper has been demonstrated to effectively improve the performance of transformers by conducting experiments on the CIFAR-10 and CIFAR-100 datasets. The current implementation in this paper serves as a preliminary validation of our insights, and further evaluation experiments involving more visual models and large language models will be conducted sequentially.
Appendix
We can also learn a metric matrix of general form, which is not necessarily restricted to the diagonal matrix. For clarity and simplicity, we present here only the $2\times 2$ matrix corresponding to two-dimensional subspace.
Let,
$\mathbf{S}=\left[\begin{matrix}s_{1}&0\\[8.09998pt] 0&s_{2}\end{matrix}\right],$
where $s_{1}>0$ and $s_{2}>0$. Then, the symmetric positive definite matrices are preset as follows,
$\mathbf{M}_{m}$ $=\left[\begin{matrix}\cos{\left(m\theta\right)}&-\sin{\left(m\theta\right)}\\[8.09998pt] \sin{\left(m\theta\right)}&\;\;\;\cos{\left(m\theta\right)}\end{matrix}\right]^{\top}\left[\begin{matrix}s_{1}^{m}&0\\[8.09998pt] 0&s_{2}^{m}\end{matrix}\right]\left[\begin{matrix}\cos{\left(m\theta\right)}&-\sin{\left(m\theta\right)}\\[8.09998pt] \sin{\left(m\theta\right)}&\;\;\;\cos{\left(m\theta\right)}\end{matrix}\right]$
$=\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{m}\cdot\breve{\bf R}(m\theta),$
and,
$\mathbf{M}_{n}$ $=\left[\begin{matrix}\cos{\left(n\theta\right)}&-\sin{\left(n\theta\right)}\\[8.09998pt] \sin{\left(n\theta\right)}&\;\;\;\cos{\left(n\theta\right)}\end{matrix}\right]^{\top}\left[\begin{matrix}s_{1}^{n}&0\\[8.09998pt] 0&s_{2}^{n}\end{matrix}\right]\left[\begin{matrix}\cos{\left(n\theta\right)}&-\sin{\left(n\theta\right)}\\[8.09998pt] \sin{\left(n\theta\right)}&\;\;\;\cos{\left(n\theta\right)}\end{matrix}\right]$
$=\breve{\bf R}(-n\theta)\cdot\mathbf{S}^{n}\cdot\breve{\bf R}(n\theta).$
Then, the linear transformations can be preset as the following matrices,
$\mathbf{T}_{m}$ $=\left[\begin{matrix}\;\;\;\cos{\left(m\theta\right)}&\sin{\left(m\theta\right)}\\[8.09998pt] -\sin{\left(m\theta\right)}&\cos{\left(m\theta\right)}\end{matrix}\right]\left[\begin{matrix}s_{1}^{-\frac{m}{2}}&0\\ 0&s_{2}^{-\frac{m}{2}}\end{matrix}\right]$
$=\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{-\frac{m}{2}},$
and,
$\mathbf{T}_{n}$ $=\left[\begin{matrix}\;\;\;\cos{\left(n\theta\right)}&\sin{\left(n\theta\right)}\\[8.09998pt] -\sin{\left(n\theta\right)}&\cos{\left(n\theta\right)}\end{matrix}\right]\left[\begin{matrix}s_{1}^{-\frac{n}{2}}&0\\ 0&s_{2}^{-\frac{n}{2}}\end{matrix}\right]$
$=\breve{\bf R}(-n\theta)\cdot\mathbf{S}^{-\frac{n}{2}}.$
It is straightforward to verify that:
$\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}=\mathbf{T}_{n}^{\top}\mathbf{M}_{n}\mathbf{T}_{n}.$
and,
$\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$ $=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{m}\cdot\breve{\bf R}(m\theta)\cdot\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{-\frac{m}{2}}\cdot\mathbf{S}^{\frac{n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}$
$=\mathbf{q}_{m}^{\top}\left[\breve{\bf R}(-m\theta)\cdot\mathbf{S}^{\frac{m+n}{2}}\cdot\breve{\bf R}(n\theta)\right]\mathbf{k}_{n}.$
Then, the form of a standard inner product is,
$\langle\mathbf{S}^{\frac{m}{2}}\breve{\bf R}(m\theta)\mathbf{q}_{m},\;\mathbf{S}^{\frac{n}{2}}\breve{\bf R}(n\theta)\mathbf{k}_{n}\rangle_{\mathbf{I}}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$